OpenStudy (anonymous):

If g(t)=4t-t^2, find g(t+h)-g(t)/h. Can someone please tell or show me step by step how to do this problem. I am having a hard time starting it

OpenStudy (anonymous):

g(t+h) means you replace all t in g(t) with t-h so it's+ g(t+h) = 4(t+h) - (t+h)^2 so g(t+h) = 4t + 4h - (t^2 + 2th + h^2) do you know how to do the rest?

OpenStudy (anonymous):

Here is the equation If g(t) = 4t - t^2, Find g(t +h) - g(t)/h

OpenStudy (accessdenied):

To find the full equation, you have to know g(t+h) first! So you see @xlegendx 's post above is telling you how to obtain g(t+h). Having both g(t+h) and g(t) then, you can compute g(t+h) - g(t) and divide by h.